# Examples of topologically unknotted tori

### v2

@@ 5. Knot surgery
be the result of $t$-twist $r$-roll knot surgery along $T$ with pattern $J$, where we describe the gluing map $f$ next.
@@
Since $T$ is smoothly unknotted in $S^{4}$, there is a diffeomorphism $\Theta\colon S^{4}_{t,r}(T,J)\to S^{4}$; see Kim–Ruberman [KR08]. We write
@@
Our discussion below works for arbitrary $g$, but it follows from the work of Conway and the second author [CP23] that $\Sigma_{t,r}(T,J)$ is topologically standard if it is a $\mathbb{Z}$-surface and $g\not\in\{1,2\}$. Let $\boldsymbol{\alpha}=\{\alpha_{1},\dots,\alpha_{g}\}$ and $\boldsymbol{\beta}=\{\beta_{1},\dots,\beta_{g}\}$ be collections of pairwise disjoint simple closed curves on $\Sigma$ such that $\boldsymbol{\alpha}\cup\boldsymbol{\beta}$ represents a basis of $H_{1}(\Sigma)$. Furthermore, we can assume that $|\alpha_{i}\pitchfork\beta_{j}|=\delta_{ij}$.
Let $R_{\gamma}\subseteq S^{4}\setminus\Sigma$ be a rim torus around $\gamma$ for $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$. Let $S_{\gamma}\subseteq S^{4}\setminus\Sigma$ be a topologically embedded 2-sphere obtained by surgering $R_{\gamma}$ along a topologically embedded disc. This is possible since $\Sigma$ is topologically unknotted, and so a longitude of $R_{\gamma}$ that links $\Sigma$ trivially bounds a topologically embedded disc in $S^{4}\setminus N(\Sigma)$. We may and shall choose these discs to be pairwise disjoint. Furthermore, we choose the disc with boundary on $R_{\alpha_{i}}$ to be disjoint from all the other rim tori except $R_{\beta_{i}}$, and the disc with boundary on $R_{\beta_{i}}$ to be disjoint from all the other rim tori except $R_{\alpha_{i}}$. It follows that
@@
For each $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$, construct $S_{\gamma}^{\prime}\subseteq S^{4}_{t,r}(T,J)$ by first making $S_{\gamma}$ transverse to $T$, and replacing each component of $S_{\gamma}\cap N(T)$, which is a disc, with a Seifert surface of $J$ in $E_{J}\times\{p\}$ for some $p\in S^{1}$. We can assume that we have a different $p$ for every component, so these Seifert surfaces are mutually disjoint.
@@
We can arrange that $S_{\alpha_{i}}\cap S_{\beta_{j}}\subseteq S^{4}\setminus N(\Sigma\cup T)$ for every $i$, $j\in\{1,\dots,g\}$. Then $S_{\alpha_{i}}^{\prime}\cap S_{\beta_{j}}^{\prime}\subseteq S^{4}\setminus N(\Sigma\cup T)$ for every $i$, $j\in\{1,\dots,g\}$ as well. Let $U_{\gamma}:=\Theta(S_{\gamma}^{\prime})\subseteq E$ for $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$, oriented in such a way that $\Theta|_{S_{\gamma}^{\prime}}\colon S_{\gamma}^{\prime}\to U_{\gamma}$ is orientation-preserving. We compute the intersection numbers among the $U_{\gamma}$. Since each $S_{\gamma}^{\prime}$ is an embedded $\mathbb{Z}$-surface, each $U_{\gamma}$ is an embedded $\mathbb{Z}$-surface. Since $U_{\gamma}\subseteq S^{4}$, it has trivial normal bundle. Thus $\lambda_{E}([U_{\gamma}],[U_{\gamma}])=0$ for every $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$. Since $\Theta$ is a diffeomorphism, the analogous facts in (6) for the $S_{\gamma}^{\prime}$ imply that
@@
Let $q\in S_{\alpha_{i}}\cap S_{\beta_{i}}$ be an intersection point with sign $\varepsilon\in\{+,-\}$ and double point group element $t^{k}\in\langle t\rangle\cong C_{\infty}\cong\mathbb{Z}$. Since $q\in S^{4}\setminus N(\Sigma\cup T)$, we can consider $\Theta(q)\in U_{\alpha_{i}}\cap U_{\beta_{i}}\subseteq S^{4}\setminus N(\Sigma_{t,r}(T,J)$. Then $\Theta(q)$ also has sign $\varepsilon$ and double point group element $t^{k}$. It follows that $\lambda_{E}([U_{\alpha_{i}}],[U_{\beta_{i}}])=\lambda_{E_{\Sigma}}([S_{\alpha_{i}}],[S_{\beta_{i}}])=1-t$.
@@
As $\Theta$ is orientation-preserving and the orientations of $U_{\alpha_{i}}$ and $U_{\beta_{i}}$ are induced by $\Theta$, the intersection point $\Theta(q)$ has the same sign as $q$. Let $V\subseteq S^{4}$ be a disc such that $\partial V$ is a double point loop for $q$. Then $V$ intersects $\Sigma$ transversely in $k$ points, counted with sign. We can arrange that $V$ is transverse to $T$. Replace $V\cap N(T)$ with Seifert surfaces of $J$ to obtain $V^{\prime}\subseteq S^{4}_{t,r}(T,J)$, and write $W:=\Theta(V^{\prime})\subseteq S^{4}$. Then $\partial W$ is a double point loop for $\Theta(q)$, and $W$ intersects $\Sigma_{t,r}(T,J)$ transversely in $k$ points, again counted with sign. As $\Theta\colon S^{4}_{t,r}(T,J)\setminus N(\Sigma)\to S^{4}\setminus N(\Sigma_{t,r}(T,J))$ induces an isomorphism on fundamental groups, both of which are isomorphic to $\mathbb{Z}$, it follows that the double point group element of $\Theta(q)$ is $t^{k}$. We deduce that $\lambda_{E}([U_{\alpha_{i}}],[U_{\beta_{i}}])=\lambda_{E_{\Sigma}}([S_{\alpha_{i}}],[S_{\beta_{i}}])$ for $i\in\{1,\dots,g\}$. ∎
@@ 6. The Cochran–Davis Seifert surface union a ribbon disc
A natural idea for constructing potentially nontrivial, closed, orientable $\mathbb{Z}$-surfaces, smoothly embedded in $S^{4}$, is to take the union of a Seifert surface $F$ for a slice knot $K$ in $S^{3}$ and a homotopy ribbon disc $D$ for $K$. However, by the following lemma, this is smoothly unknotted if $D$ is obtained from $F$ by compressing along a slice derivative, which is often the case.
@@
Since $D$ is obtained by surgering $F$ along $c$, it follows that $\Sigma$ bounds an embedded copy of $F\times I$ with a collection of 3-dimensional 2-handles attached along $c\times\{0\}$, which we denote by $H$. Since the components of $c$ are $g$ homologically linearly independent simple closed curves in $F$, the 3-manifold $H$ is a genus $g$ handlebody. As $\partial H=\Sigma$, it follows that $\Sigma$ is smoothly unknotted. ∎
Kauffman conjectured that every genus one Seifert surface for a slice knot admits a slice derivative, but Cochran and Davis [CD15] constructed a counterexample. This motivates 1.2, which we restate for the reader’s convenience. See 1.2
@@
Hence $\Sigma$ is obtained from $\Sigma^{\prime}=F^{\prime}\cup\Delta$ via knot surgery along the torus $T_{A}:=A_{+}\cup A_{-}$ with pattern the knot $J$.
Note that $\pi_{1}(D^{4}\setminus F^{\prime})\cong\mathbb{Z}$, generated by a meridian by Chen [Che21, Lemma 3.1]. Since $\Delta$ is homotopy ribbon, it follows from the Seifert–van Kampen theorem that $\Sigma^{\prime}$ is a $\mathbb{Z}$-torus.
@@ 7. Tori with four critical points
By 2.2, we already knew that $H_{2}(E_{\Sigma};\Lambda)\cong\Lambda\oplus\Lambda$, but this short exact sequence gives us a basis $\{[S_{0}],[S_{1}]\}$, as follows. Let $S_{0}\subseteq E_{L}$ denote a surface generating $H_{2}(E_{L};\Lambda)$. The intersection number $\lambda([S_{0}],[S_{0}])$ is trivial, because we can use the boundary collar $E_{L}\times[0,1]\subseteq W_{-}$ to obtain a disjoint copy of $S_{0}$ that represents the same class in $H_{2}(E_{\Sigma};\Lambda)$.
@@
Consider a standard, oriented, 2-component unlink $L$ and a band $B$ with one end attached to each component of $L$, and suppose that $L\cup B$ is isotopic to the standard oriented unlink $L$ with a standard band $B^{\prime}$ (normal-(((see the top of Figure 5)normal-))). Then $B$ is isotopic to $B^{\prime}$ relative to its ends, via an isotopy that fixes $L$ setwise throughout.
@@ 8. Double branched covers
disjoint from $K$. Then $S_{0}$ lifts to a pair of spheres in $\Sigma(S^{4},K)$, one of which we denote by $S$. Since $S_{0}$ has self-intersection 00, so does $S$. The disc $D$ lifts to a sphere $S^{\prime}$ in $\Sigma(S^{4},K)$, which has even self-intersection (in fact, it is $4$ by the proof of Theorem 1.1). Furthermore, $|S\pitchfork S^{\prime}|=1$, so $N(S\cup S^{\prime})$ is a plumbing of two spheres of self-intersection 0 and 4, respectively, and is hence diffeomorphic to $(S^{2}\times S^{2})\setminus D^{4}$; see Neumann–Weintraub [NW78]. As $T_{K,1,1}$ and $T^{2}$ are topologically isotopic by Theorem 1.1 and $\Sigma(S^{4},T^{2})\cong S^{2}\times S^{2}$, it follows that $\Sigma(S^{4},T_{K,1,1})$ is homeomorphic to $S^{2}\times S^{2}$. Consequently,
@@
We have the following consequence of Proposition 8.1.
@@
For a 4-manifold $M$ with $b_{2}^{+}(M)=1$ and $b_{2}^{-}(M)<10$, the Seiberg–Witten invariant is well-defined in the metric chamber; see Szabó [Sza96, Lemma 3.2]. In particular, if $M$ is $(S^{2}\times S^{2})\#X$ for a homotopy 4-sphere $X$, then, for every characteristic element $K\in H^{2}(X)$, Riemannian metric $g$, and small enough perturbing 2-form $h$, the Seiberg–Witten invariant $\text{SW}_{M}(K,g,h)$ is well-defined, and it vanishes by the adjunction inequality applied to the self-intersection zero 2-sphere $S^{2}\times\{p\}$ for $p\in S^{2}$ in $M$. Hence, the Seiberg–Witten invariants of $\Sigma(S^{4},T_{K,1,1})$ and $S^{2}\times S^{2}$ agree in all chambers by the wall-crossing formula. ∎
In particular, we cannot use the Seiberg–Witten invariants of the double branched covers to distinguish $T_{K,1,1}$ from $T^{2}$ up to diffeomorphism of pairs. However, this might still be a possibility if $T$ has four critical points and $T\cap S^{3}$ is not split, but we do not have any such examples that are not already known to be smoothly standard. We finally consider the double cover $M$ of $S^{4}$ branched along the Cochran–Davis Seifert surface union a ribbon disc. By Theorem 1.2, the 4-manifold $M$ is homeomorphic to $S^{2}\times S^{2}$. On the other hand, we have the following.
@@
Let $p\colon M\to S^{2}\times S^{2}$ be the covering map. We now show that $T:=p^{-1}(T_{A})$ is connected, and hence a torus. The curve $\eta_{1}$ is a meridian of $T_{A}$. Since $\text{lk}(\eta_{1},R)=0$ and $\Sigma^{\prime}\cap S^{3}=R$, the meridian of $T_{A}$ is 0-homologous in $S^{4}\setminus\Sigma^{\prime}$. We claim that a longitude of $T_{A}$ represents a generator of $H_{1}(S^{4}\setminus\Sigma^{\prime})\cong\mathbb{Z}$. Recall that the annulus $A_{\pm}$ is obtained by performing a saddle move on $\eta_{1}\cup\eta_{2}$ and capping off the resulting unknot with a disc. Suppose that the saddle is attached at points $x_{1}\in\eta_{1}$ and $x_{2}\in\eta_{2}$; see Figure 6. We can assume that $x_{1}$ and $x_{2}$ lie on opposite sides of the Seifert surface $F^{\prime}$. Suppose that the saddle point of $A_{\pm}$ lies in $S^{3}\times\{t_{\pm}\}$, and let $c_{\pm}$ be the core of the band that is attached. Then
@@
Let $a$ be a straight arc in $S^{3}$ connecting $x_{1}$ and $x_{2}$ such that $|a\pitchfork F^{\prime}|=1$; see Figure 6. Then $D_{A}:=a\times[t_{-},t_{+}]$ gives a disc in $S^{4}$ with boundary $\ell$. We have $D_{A}\pitchfork\digamma^{\prime}=\{(a\cap F^{\prime},-1/8)\}$. Furthermore, $D_{A}\cap\Delta=\emptyset$, because we attach the saddle to $\eta_{1}\cup\eta_{2}$ to obtain $A_{+}$ before we attach the saddle to $R$ when forming $\Delta$. It follows that $|D_{A}\pitchfork\Sigma^{\prime}|=1$. Hence $\ell$ represents a generator of $H_{1}(S^{4}\setminus\Sigma^{\prime})$, as claimed. Consequently, the pre-image $p^{-1}(T_{A})$ is connected. By construction, we obtain $\Sigma$ from $\Sigma^{\prime}$ by performing knot surgery along $T_{A}$ with pattern some knot $J$ in $S^{3}$ (in the main Cochran–Davis example, $J$ is the left-handed trefoil). So we can obtain $M$ from $S^{2}\times S^{2}$ by knot surgery along the torus $T$. Since $H_{2}(S^{4})=0$, the torus $T_{A}$ bounds a 3-manifold $N$ in $S^{4}$. Make $N$ transverse to $\Sigma^{\prime}$. Then $p^{-1}(N)$ is a 3-manifold in $M$ with boundary $T$. Hence $T$ is 0-homologous, as claimed. ∎

### v3

@@ 5. Knot surgery
be the result of $t$-twist $r$-roll knot surgery along $T$ with pattern $J$, where we describe the glueing map $f$ next.
@@
Since $T$ is smoothly unknotted in $S^{4}$, there is a diffeomorphism
$\Theta\colon S^{4}_{t,r}(T,J)\to S^{4};$
see Kim–Ruberman [KR08]. We write
@@
Our discussion below works for arbitrary $g$, but it follows from the work of Conway and the second author [CP23] that $\Sigma_{t,r}(T,J)$ is topologically standard if it is a $\mathbb{Z}$-surface and $g\not\in\{1,2\}$. The strategy is as follows. Let $E:=S^{4}\setminus N(\Sigma_{t,r}(T,J))$ and $E_{\Sigma}:=S^{4}\setminus N(\Sigma)$. We describe generators for $H_{2}(E_{\Sigma};\Lambda)$, and then modify them during the knot surgery construction to obtain generators for $H_{2}(E;\Lambda)$. We check that the intersection form is unchanged.
Now we begin to implement this. Since $\Sigma$ is topologically unknotted, there is a locally flat 3-manifold $Y\subseteq S^{4}$ homeomorphic to $S^{3}$ such that $\Sigma\subseteq Y$ and the closures $H_{\alpha}$ and $H_{\beta}$ of the components of $Y\setminus\Sigma$ are handlebodies. Let $\boldsymbol{\alpha}=\{\alpha_{1},\dots,\alpha_{g}\}$ and $\boldsymbol{\beta}=\{\beta_{1},\dots,\beta_{g}\}$ be collections of pairwise disjoint simple closed curves on $\Sigma$ such that $\boldsymbol{\alpha}\cup\boldsymbol{\beta}$ represents a basis of $H_{1}(\Sigma)$, and $\alpha_{i}$ bounds a locally flat disc $D_{\alpha_{i}}\subseteq H_{\alpha}$ and $\beta_{i}$ bounds a locally flat disc $D_{\beta_{i}}\subseteq H_{\beta}$ for $i\in\{1,\dots,g\}$. Furthermore, we can assume that $|\alpha_{i}\pitchfork\beta_{j}|=\delta_{ij}$ and the interiors of the discs $D_{\gamma}$ for $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$ are pairwise disjoint.
Let $R_{\gamma}\subseteq S^{4}\setminus\Sigma$ be a rim torus around $\gamma$ for $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$, which is the boundary of the restriction $N_{\gamma}$ of the unit normal disc bundle of $\Sigma$ to $\gamma$. Let $S_{\gamma}\subseteq S^{4}\setminus\Sigma$ be the topologically embedded 2-sphere obtained by surgering $R_{\gamma}$ along the topologically embedded disc $D_{\gamma}^{\prime}:=D_{\gamma}\setminus\operatorname{Int}(N_{\gamma})$. We can arrange that $D_{\alpha_{i}}^{\prime}\cap R_{\gamma}=\emptyset$ and $D_{\beta_{i}}^{\prime}\cap R_{\gamma}=\emptyset$ for any $i\in\{1,\dots,g\}$ and $\gamma\in(\boldsymbol{\alpha}\cup\boldsymbol{\beta})\setminus\{\alpha_{i},\beta_{i}\}$. It follows that
@@
The intersection $S_{\gamma}\cap N(T)$ is a collection of meridional discs. The boundary of each disc is a meridian of $T$. In the definition of the glueing map $f\colon\partial\overline{N(T)}\to\partial E_{J}\times S^{1}$, we have that $f(\mu_{T})=\lambda_{J}\times\{1\}$. Hence each meridian of $T$ is taken to a parallel of $J$. For each $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$, construct $S_{\gamma}^{\prime}\subseteq S^{4}_{t,r}(T,J)$ by first making $S_{\gamma}$ transverse to $T$, and replacing each component of $S_{\gamma}\cap N(T)$, which is a disc, with a Seifert surface of $J$ in $E_{J}\times\{p\}$ for some $p\in S^{1}$. We can assume that we have a different $p$ for every component, so these Seifert surfaces are mutually disjoint.
@@
We can arrange that $S_{\alpha_{i}}\cap S_{\beta_{j}}\subseteq S^{4}\setminus N(\Sigma\cup T)$ for every $i$, $j\in\{1,\dots,g\}$. Then $S_{\alpha_{i}}^{\prime}\cap S_{\beta_{j}}^{\prime}\subseteq S^{4}\setminus N(\Sigma\cup T)$ for every $i$, $j\in\{1,\dots,g\}$ as well. Let
$U_{\gamma}:=\Theta(S_{\gamma}^{\prime})\subseteq E$
for $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$, oriented in such a way that $\Theta|_{S_{\gamma}^{\prime}}\colon S_{\gamma}^{\prime}\to U_{\gamma}$ is orientation-preserving. We compute the intersection numbers among the $U_{\gamma}$. Since each $S_{\gamma}^{\prime}$ is an embedded $\mathbb{Z}$-surface, each $U_{\gamma}$ is an embedded $\mathbb{Z}$-surface. Since $U_{\gamma}\subseteq S^{4}$, it has trivial normal bundle. Thus $\lambda_{E}([U_{\gamma}],[U_{\gamma}])=0$ for every $\gamma\in\boldsymbol{\alpha}\cup\boldsymbol{\beta}$. Since $\Theta$ is a diffeomorphism, the analogous facts in (6) for the $S_{\gamma}^{\prime}$ imply that
@@
Let $q\in S_{\alpha_{i}}\cap S_{\beta_{i}}$ be an intersection point with sign $\varepsilon\in\{+,-\}$ and double point group element $t^{k}\in\langle t\rangle\cong C_{\infty}\cong\mathbb{Z}$. Since $q\in S^{4}\setminus N(\Sigma\cup T)$, we can consider $\Theta(q)\in U_{\alpha_{i}}\cap U_{\beta_{i}}\subseteq S^{4}\setminus N(\Sigma_{t,r}(T,J))$. Then $\Theta(q)$ also has sign $\varepsilon$ and double point group element $t^{k}$. It follows that $\lambda_{E}([U_{\alpha_{i}}],[U_{\beta_{i}}])=\lambda_{E_{\Sigma}}([S_{\alpha_{i}}],[S_{\beta_{i}}])=1-t$.
@@
As $\Theta$ is orientation-preserving and the orientations of $U_{\alpha_{i}}$ and $U_{\beta_{i}}$ are induced by $\Theta$, the intersection point $\Theta(q)$ has the same sign as $q$.
Let $V\subseteq S^{4}$ be a disc such that $\partial V$ is a double point loop for $q$. Then $V$ intersects $\Sigma$ transversely in $k$ points, counted with sign. We can arrange that $V$ is transverse to $T$. Replace $V\cap N(T)$ with Seifert surfaces of $J$ to obtain $V^{\prime}\subseteq S^{4}_{t,r}(T,J)$, and write $W:=\Theta(V^{\prime})\subseteq S^{4}$. Then $\partial W$ is a double point loop for $\Theta(q)$, and $W$ intersects $\Sigma_{t,r}(T,J)$ transversely in $k$ points, again counted with sign. As $\Theta\colon S^{4}_{t,r}(T,J)\setminus N(\Sigma)\to S^{4}\setminus N(\Sigma_{t,r}(T,J))$ induces an isomorphism on fundamental groups, both of which are isomorphic to $\mathbb{Z}$, it follows that the double point group element of $\Theta(q)$ is $t^{k}$. We deduce that $\lambda_{E}([U_{\alpha_{i}}],[U_{\beta_{i}}])=\lambda_{E_{\Sigma}}([S_{\alpha_{i}}],[S_{\beta_{i}}])$ for $i\in\{1,\dots,g\}$. ∎
@@ 6. The Cochran–Davis Seifert surface union a ribbon disc
A natural idea for constructing potentially nontrivial, closed, orientable $\mathbb{Z}$-surfaces, smoothly embedded in $S^{4}$, is to take the union of a Seifert surface $F$ for a slice knot $K$ in $S^{3}$ and a homotopy ribbon disc $D$ for $K$. However, by the following lemma, this is smoothly unknotted if $D$ is obtained from $F$ by compressing along a slice derivative, which is often the case. Recall that a slice derivative for a genus $g$ Seifert surface $F$ for a knot $K$ in $S^{3}$ is a collection of $g$ pairwise disjoint, homologically linearly independent simple closed curves that form a smoothly slice link such that the Seifert form vanishes identically on the submodule of $H_{1}(F)$ generated by these curves.
@@
Since $D$ is obtained by surgering $F$ along $c$, it follows that $\Sigma$ bounds an embedded copy of $F\times I$ with a collection of 3-dimensional 2-handles attached along $c\times\{0\}$, which we denote by $H$. Since the components of $c$ are $g$ homologically linearly independent simple closed curves in $F$, the 3-manifold $H$ is a genus $g$ handlebody. To see this, note that, if we turn the relative handle decomposition of $H$ built on $F\times I$ upside down, we obtain a handle decomposition of $H$ with only 1-handles relative to $D^{2}\times I$, as $D^{2}$ is the surface obtained by compressing $F$ along $c$. So one can obtain $H$ by attaching 3-dimensional 1-handles to $D^{2}\times I$, which is a handlebody. Since $\partial H=\Sigma$, it follows that $\Sigma$ is smoothly unknotted. ∎
Kauffman conjectured that every genus one Seifert surface for a slice knot admits a slice derivative, but Cochran and Davis [CD15] constructed a counterexample. This motivates 1.2, which we restate for the reader’s convenience.
See 1.2
@@
Hence $\Sigma$ is obtained from $\Sigma^{\prime}=F^{\prime}\cup\Delta$ via knot surgery along the torus $T_{A}:=A_{+}\cup A_{-}$ with pattern the knot $J$. To see this, note that knot surgery removes $N(T)$ and glues in $E_{J}\times S^{1}$. Our construction removes $N(A_{+})\cup N(A_{-})$ and glues in $(E_{J}\times[0,1])\cup(E_{J}\times[0,1])\cong E_{J}\times S^{1}$. As the glueing maps coincide, $\Sigma$ is indeed obtained via knot surgery.
Note that $\pi_{1}(D^{4}\setminus F^{\prime})\cong\mathbb{Z}$, generated by a meridian by [Che21, Lemma 3.1]. Since $\Delta$ is homotopy ribbon, it follows from the Seifert–van Kampen theorem that $\Sigma^{\prime}$ is a $\mathbb{Z}$-torus.
@@ 7. Tori with four critical points
The previous proof is the key place where we used the assumption that $\Delta_{L}=0$.
@@
By 2.2, we already knew that $H_{2}(E_{\Sigma};\Lambda)\cong\Lambda\oplus\Lambda$, but this short exact sequence gives us a basis $\{[S_{0}],[S_{1}]\}$, as follows. Let $S_{0}\subseteq E_{L}$ denote a surface generating $H_{2}(E_{L};\Lambda)$. Let $\lambda$ be the $\Lambda$-intersection form on $H_{2}(E_{\Sigma};\Lambda)$. The intersection number $\lambda([S_{0}],[S_{0}])$ is trivial, because we can use the boundary collar $E_{L}\times[0,1]\subseteq W_{-}$ to obtain a disjoint copy of $S_{0}$ that represents the same class in $H_{2}(E_{\Sigma};\Lambda)$.
@@
Consider a standard, oriented, 2-component unlink $L$ and a band $B$ with one end attached to each component of $L$, and suppose that $L\cup B$ is isotopic to the standard oriented unlink $L$ with a standard band $B^{\prime}$ ((((see the top of Figure 5)))). Then $B$ is isotopic to $B^{\prime}$ relative to its ends, via an isotopy that fixes $L$ setwise throughout.
@@ 8. Double branched covers
disjoint from $K\times\{0\}$. Then $S_{0}$ is disjoint from the branch locus $T_{K,1,1}$ and is simply connected, so it lifts to a pair of spheres in $\Sigma(S^{4},T_{K,1,1})$, one of which we denote by $S$. Since $S_{0}$ has self-intersection zero, so does $S$. As $D\cap T_{K,1,1}=\partial D$, the disc $D$ lifts to a sphere $S^{\prime}$ in $\Sigma(S^{4},T_{K,1,1})$, which has even self-intersection (in fact, it is four by the proof of Theorem 1.1). Note that $\partial D^{3}\cap J=\{(0,0,1)\}$, so $|S_{0}\cap D|=1$. It follows that $|S\pitchfork S^{\prime}|=1$, so $N(S\cup S^{\prime})$ is a plumbing of two spheres of self-intersection zero and four, respectively, and is hence diffeomorphic to $(S^{2}\times S^{2})\setminus D^{4}$; see Neumann–Weintraub [NW78].
As $T_{K,1,1}$ and $T^{2}$ are topologically isotopic by Theorem 1.1 and $\Sigma(S^{4},T^{2})\cong S^{2}\times S^{2}$, it follows that $\Sigma(S^{4},T_{K,1,1})$ is homeomorphic to $S^{2}\times S^{2}$. Consequently,
@@
Let $M$ be a closed, connected, smooth, and oriented 4-manifold $M$ with $b_{1}(M)=0$ and $b_{2}^{+}(M)>1$. Furthermore, let $o$ be an orientation of $H^{2}_{+}(M;\mathbb{R})$, which is called a homology orientation. The Seiberg–Witten invariant of $M$ is a map
$\text{SW}_{M,o}\colon\text{Spin}^{c}(M)\to\mathbb{Z},$
where $\text{Spin}^{c}(M)$ is the set of $\text{Spin}^{c}$ structures on $X$, which is an affine space modelled on $H^{2}(M)$. This satisfies $\text{SW}_{M,-o}=-\text{SW}_{M,o}$. If $\Phi\colon M^{\prime}\to M$ is a diffeomorphism, $o$ is a homology orientation of $M$, and $\mathfrak{s}\in\text{Spin}^{c}(M)$, then
$\text{SW}_{M^{\prime},\Phi^{*}o}(\Phi^{*}\mathfrak{s})=\text{SW}_{M,o}(\mathfrak{s}).$
We say that the Seiberg–Witten invariants of $M$ and $M^{\prime}$ agree if there is an affine isomorphism $\phi\colon\text{Spin}^{c}(M)\to\text{Spin}^{c}(M^{\prime})$ such that
$\text{SW}_{M^{\prime},o^{\prime}}\circ\phi=\pm\text{SW}_{M,o}$
for arbitrary homology orientations $o$ of $M$ and $o^{\prime}$ of $M^{\prime}$.
By Kronheimer and Mrowka [KM94] (see also Gompf–Stipsicz [GS99, Theorem 2.4.8]), if $\Sigma$ is a smoothly embedded, connected, and oriented surface of genus $g$ in $M$ with self-intersection $[\Sigma]^{2}\geq 0$ and $[\Sigma]\neq 0$ in $H_{2}(M)$, then
(9) $2g-2\geq[\Sigma]^{2}+|\langle c_{1}(\mathfrak{s}),[\Sigma]\rangle|$
whenever $\text{SW}_{M,o}(\mathfrak{s})\neq 0$. This is known as the generalised andjunction formula.
One can also define $\text{SW}_{M,o}(\mathfrak{s})\in\mathbb{Z}$ when $b_{1}(X)=0$ and $b_{2}^{+}(M)=1$, but it also depends on a choice of two possible metric chambers $c_{+}$ and $c_{-}$. When the expected dimension of the Seiberg–Witten moduli space
$d(\mathfrak{s}):=\left(c_{1}(\mathfrak{s})^{2}-2\chi(M)-3\sigma(M)\right)/4\in 2\mathbb{Z}_{\geq 0},$
then the dependence of the Seiberg–Witten invariant on the choice of chamber is described by the wall-crossing formula
(10) $\text{SW}_{M,o}(\mathfrak{s},c_{-})=\text{SW}_{M,o}(\mathfrak{s},c_{+})+(-1)^{d(\mathfrak{s})/2}$
due to Kronheimer and Mrowka [KM94]. We have the following consequence of Proposition 8.1.
@@
For a closed, connected, smooth, and oriented 4-manifold $M$ with $b_{2}^{+}(M)=1$ and $b_{2}^{-}(M)<10$, there is a canonical metric chamber for the Seiberg–Witten invariant; see Szabó [Sza96, Lemma 3.2]. Furthermore, the generalised adjunction formula holds with this choice. Hence, if $M$ is $(S^{2}\times S^{2})\#X$ for a homotopy 4-sphere $X$, then, for every $\mathfrak{s}\in\text{Spin}^{c}(M)$, we have $\text{SW}_{M,o}(\mathfrak{s},c)=0$ for the canonical metric chamber $c$ by the generalised adjunction formula (9) applied to the self-intersection zero, homologically non-trivial 2-sphere $\Sigma:=S^{2}\times\{p\}\subseteq M$ for $p\in S^{2}$ (as the inequality $-2\geq|\langle c_{1}(\mathfrak{s}),[\Sigma]\rangle|$ does not hold for any $\mathfrak{s}$). Hence, the Seiberg–Witten invariants of $\Sigma(S^{4},T_{K,1,1})$ and $S^{2}\times S^{2}$ both vanish in the canonical metric chambers, and also agree in the other chamber by the wall-crossing formula (10). ∎
In particular, we cannot use the Seiberg–Witten invariants of the double branched covers to distinguish $T_{K,1,1}$ from $T^{2}$ up to diffeomorphism of pairs. However, this might still be a possibility if $T$ has four critical points and $T\cap S^{3}$ is not split, but we do not have any such examples that are not already known to be smoothly standard.
We finally consider the double cover $M$ of $S^{4}$ branched along the Cochran–Davis Seifert surface union a ribbon disc. By Theorem 1.2, the 4-manifold $M$ is homeomorphic to $S^{2}\times S^{2}$. On the other hand, we have the following.
@@
Let $p\colon M\to S^{2}\times S^{2}$ be the covering map. We now show that $T:=p^{-1}(T_{A})$ is connected, and hence a torus. The curve $\eta_{1}$ is a meridian of $T_{A}$. Since $\text{lk}(\eta_{1},R)=0$ and $\Sigma^{\prime}\cap S^{3}=R$, the meridian of $T_{A}$ is 0-homologous in $S^{4}\setminus\Sigma^{\prime}$.
We claim that a longitude of $T_{A}$ represents a generator of $H_{1}(S^{4}\setminus\Sigma^{\prime})\cong\mathbb{Z}$. Recall that the annulus $A_{\pm}$ is obtained by performing a saddle move on $\eta_{1}\cup\eta_{2}$ and capping off the resulting unknot with a disc. Suppose that the saddle is attached at points $x_{1}\in\eta_{1}$ and $x_{2}\in\eta_{2}$; see Figure 6. We can assume that $x_{1}$ and $x_{2}$ lie on opposite sides of the Seifert surface $F^{\prime}$. Suppose that the saddle point of $A_{\pm}$ lies in $S^{3}\times\{t_{\pm}\}$, and let $c_{\pm}$ be the core of the band that is attached. Then
@@
Let $a$ be a straight arc in $S^{3}$ connecting $x_{1}$ and $x_{2}$ such that $|a\pitchfork F^{\prime}|=1$; see Figure 6. Then $D_{A}:=a\times[t_{-},t_{+}]$ gives a disc in $S^{4}$ with boundary $\ell$. We have $D_{A}\pitchfork\digamma^{\prime}=\{(a\cap F^{\prime},-1/8)\}$. Furthermore, $D_{A}\cap\Delta=\emptyset$, because we attach the saddle to $\eta_{1}\cup\eta_{2}$ to obtain $A_{+}$ before we attach the saddle to $R$ when forming $\Delta$. It follows that $|D_{A}\pitchfork\Sigma^{\prime}|=1$. Hence $\ell$ represents a generator of $H_{1}(S^{4}\setminus\Sigma^{\prime})$, as claimed. Consequently, the pre-image $p^{-1}(T_{A})$ is connected.
By construction, we obtain $\Sigma$ from $\Sigma^{\prime}$ by performing knot surgery along $T_{A}$ with pattern some knot $J$ in $S^{3}$ (in the main Cochran–Davis example, $J$ is the left-handed trefoil). So we can obtain $M$ from $S^{2}\times S^{2}$ by knot surgery along the torus $T$. Since $H_{2}(S^{4})=0$, the torus $T_{A}$ bounds a 3-manifold $N$ in $S^{4}$. Make $N$ transverse to $\Sigma^{\prime}$. Then $p^{-1}(N)$ is a 3-manifold in $M$ with boundary $T$. Hence $T$ is 0-homologous, as claimed. ∎